### v1

# Projection Theorems in the Presence of Expansions

Abstract.
We prove a restricted projection theorem for a certain one dimensional family of projections from $\mathbb{R}^{n}$ to $\mathbb{R}^{k}$.
The family we consider here arises naturally in the study of quantitative equidistribution problems in homogeneous dynamics.
1. Introduction
Restricted projection problems are intimately related to central questions in Fourier analysis and incidence geometry, and have been much studied, e.g., by Mattila, Falconer, Bourgain and others. More recently, certain restricted projection theorems have also found striking applications in homogeneous dynamics.
Let us recall the the classical Marstrand projection theorem: Let $K\subset\mathbb{R}^{n}$ be a compact subset, then for a.e. $v\in\mathbb{S}^{n-1}$
(1.1) $\dim{\rm p}_{v}(K)=\min(1,\dim K),$
where ${\rm p}_{v}(w)=w\cdot v$ is the orthogonal projection in the direction of $v$ and here and in what follows $\dim$ denotes the Hausdorff dimension. Analogous statements hold more generally for orthogonal projection into a.e. $m$-dimensional subspace, with respect to the Lebesgue measure on ${\rm Gr}(m,n)$.
Broadly speaking, restricted projection problems seek to obtain similar results as in (1.1) where $v$ is confined to a proper Borel subset $\mathsf{B}\subset\mathbb{S}^{n-1}$. Note, however, that without further restrictions on $\mathsf{B}$, (1.1) fails: e.g., if
$\mathsf{B}=\{(\cos t,\sin t,0):0\leq t\leq 2\pi\}$
is the great circle in $\mathbb{S}^{2}$ and $K$ is the $z$-axis, then ${\rm p}_{v}(K)=0$ for every $v\in\mathsf{B}$.
To avoid degenerations of this nature, one may, e.g., consider a curve $\gamma:[0,1]\to\mathbb{R}^{n}$ satisfying that
$\{\gamma^{\prime}(r),\gamma^{\prime\prime}(r),\ldots,\gamma^{(n)}(r)\}\qquad\text{spans $\mathbb{R}^{n}$ for all $r\in[0,1]$.}$
It is natural to inquire whether for almost all $r\in[0,1]$ the following holds
(1.2) $\dim{\rm p}_{\gamma^{\prime}(r)}(K)=\min(1,\dim K);$
again $K\subset\mathbb{R}^{n}$ is a compact subset.
Indeed (1.2) was conjectured by Fässler and Orponen [FO14] in dimension $3$. This conjecture was resolved by Käenmäki, Orponen, and Venieri [KOV17] and Pramanik, Yang, and Zahl [PYZ22] — these works rely on the work of Wolff and Schlag on circular Kakeya sets [Wol00, Sch03].
More recently, Gan, Guo, and Wang [GGW22] have established (1.2) in all dimensions using decoupling inequalities for the moment curve due to Bourgain, Demeter, and Guth [BDG16].
In this paper, we study a closely related family of projections. The consideration of these families are mainly motivated by the aforementioned applications to homogeneous dynamics.
Let us fix some notation in order to state the main results of this paper. For every $t\geq 0$ and $r\in[0,1]$, let $\pi_{t,r}:\mathbb{R}^{3}\to\mathbb{R}^{2}$ be
(1.3) $\pi_{t,r}(x,y,z)=\Bigl{(}e^{t}(x+ry+\tfrac{r^{2}}{2}z),y+rz\Bigr{)}$
The following is the main theorem in this paper.
1.1 Theorem.
Let $21\leq\alpha\leq 3/$ 2, and let $0<\delta_{0}\leq 1$. Let $F\subset B_{\mathbb{R}^{3}}(0,1)$ be a finite set satisfying the following:
#$\#(B_{\mathbb{R}^{3}}(w,\delta)\cap F)\leq C_{0}\cdot\delta^{\alpha}\cdot(\#F)\quad\text{for all $w\in\mathbb{R}^{3}$ and all $\delta\geq\delta_{0}$}$
where $C_{0}\geq 1$.
Let $1000<\varepsilon<\alpha/$ 100. For every $\delta\geq e^{t}\delta_{0}$, there is a subset $I_{\delta}\subset[1/2,1]$ with
$|[1/2,1]\setminus I_{\delta}|\ll_{\varepsilon}\delta^{\star\varepsilon}$
so that the following holds. Let $r\in I_{\delta}$, then there exists $F_{\delta,r}\subset F$ with
#$\#(F\setminus F_{\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon}\cdot(\#F)$
such that for all $w\in F_{\delta,r}$, we have
#($\#\Bigl{(}\{w^{\prime}\in F_{\delta,r}:\|\pi_{t,r}(w^{\prime})-\pi_{t,r}(w)|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}e^{-t/10}\delta^{\alpha-\varepsilon}\cdot(\#F)$
Remark.
Throughout the paper, the notation $a\ll b$ and $a^{\star b}$ mean $a\leq Db$ and $a^{Db}$, respectively, where $D$ is some positive constant whose dependence is explicated in different statements. Also, for a Borel subset $\mathsf{B}\subset\mathbb{R}^{d}$, we denote the Lebesgue measure of $\mathsf{B}$ by $|\mathsf{B}|$.
We also state a similar result in all dimensions $n\geq 3$, which relies on the aforementioned work of [GGW22].
Let $12\leq k\leq n-$ 1. For all $t\geq 0$ and all $r\in[0,1]$, let $\pi_{t,r}^{(k)}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ be
$\pi_{t,r}^{(k)}(w)=\Bigl{(}e^{t}w\cdot\xi^{\prime}(r),e^{t}w\cdot\xi^{\prime\prime}(r),\ldots,e^{t}w\cdot\xi^{(k-1)}(r),w\cdot\xi^{(k)}(r)\Bigr{)}$
where $w\cdot v$ is the usual inner product on $\mathbb{R}^{n}$ and
$\xi(r)=(\tfrac{r}{1!},\tfrac{r^{2}}{2!},\ldots,\tfrac{r^{n}}{n!})\subset\mathbb{R}^{n}.$
Using [GGW22, Thm. 2.1] and some elementary arguments, one can establish the following
1.2 Proposition.
Let $0<\alpha\leq k$, and let $0<\delta_{0}\leq 1$. Let $F\subset B_{\mathbb{R}^{n}}(0,1)$ be a finite set satisfying the following:
#$\#(B_{\mathbb{R}^{n}}(w,\delta)\cap F)\leq C_{0}\cdot\delta^{\alpha}\cdot(\#F)\quad\text{for all $w\in F$ and all $\delta\geq\delta_{0}$}$
where $C_{0}\geq 1$.
Let $1000<\varepsilon<\alpha/$ 100. For every $\delta\geq e^{t}\delta_{0}$, there is a subset $I_{\delta}\subset[1/2,1]$ with
$|[1/2,1]\setminus I_{\delta}|\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}$
so that the following holds. Let $r\in I_{\delta}$, then there exists $F_{\delta,r}\subset F$ with
#$\#(F\setminus F_{\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}\cdot(\#F)$
such that for all $w\in F_{\delta,r}$, we have
#($\#\Bigl{(}\{w^{\prime}\in F_{\delta,r}:\|\pi_{t,r}(w^{\prime})-\pi_{t,r}(w)|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}e^{-\alpha^{\prime}t}\delta^{\alpha-\varepsilon}\cdot(\#F)$
where $\alpha^{\prime}=\alpha$ if $10<\alpha\leq k-$ 1 and $\alpha^{\prime}=\alpha-1$ otherwise.
It is worth noting that Proposition 1.2 does not imply Theorem 1.1; indeed $\alpha^{\prime}=\alpha-1<1/10$ for $101\leq\alpha<11/$ 10. That said, it seems plausible that the approach developed in [GGW22] can be modified to prove a stronger version of Proposition 1.2 which then would imply (and perhaps improve) Theorem 1.1. However, our proof of Theorem 1.1 relies on much softer arguments and is an adaptation of [OV18].
Also as previously mentioned, our precise formulations here (which at least at first glance may appear somewhat ad hoc) are motivated by recent applications in homogeneous dynamics [LM23, LMW22, LMWY23].
2. Proof of Theorem 1.1
In this section we will prove Theorem 1.1. Let us begin by fixing some notation which will be used throughout this section.
Let $\mu$ denote the normalized counting measure on $F$:
$\mu(\mathsf{B})=\frac{\#(\mathsf{B}\cap F)}{\#F}\qquad\text{for any Borel set $\mathsf{B}\subset\mathbb{R}^{3}$}.$
Since $t$ is fixed throughout the argument, we will write $\pi_{r}$ for $\pi_{t,r}$. For every $w\in F$, all $r\in[0,1]$, and all $b>0$, let
$m^{b}(\pi_{r}(w))=\mu(\{w^{\prime}\in F:\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq b\}).$
More generally, given a subset $\mathsf{B}\subset\mathbb{R}^{3}$, let
$m^{b}(\pi_{r}(w)|\mathsf{B})=\mu(\{w^{\prime}\in\mathsf{B}:\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq b\}).$
For all $w\in F$ and all $b>0$, let $\mathsf{D}_{b}(w)=\{w^{\prime}:b\leq\|w-w^{\prime}\|\leq 2b\}$.
The following lemma is the main step in the proof of Theorem 1.1.
2.1 Lemma.
The following holds for all small enough $\eta$ and all large enough $C$. Let $F_{\rm hm}\subset F$ denote the set of $w\in F$ so that
|$|\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\}|\geq C\delta^{\eta}.$
Then $\mu(F_{\rm hm})\leq C\delta^{\eta}$.
We will use the following elementary lemma in the proof of Lemma 2.1.
2.2 Lemma.
Let $\sigma$ be a probability measure on $B(0,10)$ which satisfies
$\sigma(B(w,r))\leq\hat{C}r^{\beta}\quad\text{for all $w\in\mathbb{R}^{3}$ and all $r\geq r_{0}$}$
where $1/2\leq\beta\leq 1$ and $\hat{C}>0$.
The following holds for all $\varepsilon_{0}$ small enough. For every $b\geq r_{0}$, there exists a subset $E_{b}$ with $\sigma(B(0,10)\setminus E_{b})\leq b^{\varepsilon_{0}}$, and for every $z\in E_{b}$, there is a subset $I_{z}\subset[0,1]$ with $|[0,1]\setminus I_{z}|\leq b^{\varepsilon_{0}}$ so that for every $r\in I_{z}$
$\sigma\{z^{\prime}:|(1,r,\tfrac{r^{2}}{2})\cdot(z-z^{\prime})|\leq b\}\ll\hat{C}b^{49\beta/100}$
Proof.
This, rather weak estimate, follows from standard arguments, see e.g. [Ohm23, §3]. It is worth noting that much stronger result holds where $10049\beta/$ 100 is replaced by $\beta-o(1)$, see [LM23, Thm.  B.1] and [KOV17]. ∎
2.3 Lemma.
Let the notation and assumptions be as in Lemma 2.1. The following holds for all small enough $\eta$. There exists some $b\geq\delta^{1-3\eta}$ and a subset $F^{\prime}_{\rm hm}\subset F_{\rm hm}$ with
(2.1) $\mu(F^{\prime}_{\rm hm})\geq\delta^{3\eta/2}$
so that for all $w\in F^{\prime}_{\rm hm}$, we have
(2.2) |$|\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-7\eta}\}|\geq\delta^{3\eta/2}.$
Proof.
Let $b_{0}=\delta^{1-3\eta}$ and recall that $21\leq\alpha\leq 3/$ 2.
If $e^{-t}\geq\delta^{100\eta}$, then $e^{-t/10}\delta^{\alpha-18\eta}\geq\delta^{\alpha-8\eta}$. Therefore,
$\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\}\subset\\ \{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w))\geq C_{0}\delta^{\alpha-8\eta}\}.$
Now since $\mu(B(w,b_{0}))\leq C_{0}b_{0}^{\alpha}\leq C_{0}\delta^{\alpha-4.5\eta}$, there is some $b=b(w,r)\geq\delta^{1-3\eta}$ so that
(2.3) $m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-7.5\eta}.$
In view of this we assume $e^{-t}\leq\delta^{100\eta}$ for the rest of the argument. Put
$\Xi_{\delta}:=\{w:\mu(B(w,b_{0}))>C_{0}e^{-t/10}\delta^{\alpha-7\eta}\}$
Also for every $w\in F_{\rm hm}$, let
$I(w)=\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\cdot(\#F)\}.$
We now consider two possibilities:
Case 1: Assume $\mu(F_{\rm hm}\setminus\Xi_{\delta})\geq\delta^{\eta}/100$.
In this case, let $F^{\prime\prime}_{\rm hm}=F_{\rm hm}\setminus\Xi_{\delta}$. Then $\mu(F^{\prime\prime}_{\rm hm})\geq\delta^{\eta}/100$. Moreover, for every $w\in F^{\prime\prime}_{\rm hm}$, we have $\mu(B(w,b_{0}))\leq C_{0}e^{-t/10}\delta^{\alpha-7\eta}$.
Now since for every $r\in I(w)$, we have
$m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}$
there exists some $b=b(w,r)\geq\delta^{1-3\eta}$ so that
(2.4) $m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\gg C_{0}e^{-t/10}\delta^{\alpha-17\eta}.$
Case 2: Assume $\mu(F_{\rm hm}\setminus\Xi_{\delta})\leq\delta^{\eta}/100$.
In this case, we have $\mu(F_{\rm hm}\cap\Xi_{\delta})\geq\frac{C}{2}\delta^{\eta}$.
Fix a maximal $b_{0}/2$ separated subset $\{w_{1},\ldots,w_{M}\}$ of $F_{\rm hm}\cap\Xi_{\delta}$. Discarding a subset of $F_{\rm hm}$ with measure $\leq\mu(F_{\rm hm})/100$, we will assume that
$\mu(B(w_{i},b_{0}))\gg\mu(B(w_{i},4b_{0}))\qquad\text{ for all $1\leq i\leq M$},$
where the implied constant is absolute.
Let $\nu_{i}$ be the measure on $B(0,1)$ which is the image of
$\tfrac{1}{\mu(B(w_{i},b_{0}))}\mu|_{B(w_{i},b_{0})}$
under the map $w\to\frac{w_{i}-w}{b_{0}}$. Similarly, let $\tilde{\nu}_{i}$ denote the measure on $B(0,4)$ is the image of
$\tfrac{1}{\mu(B(w_{i},4b_{0}))}\mu|_{B(w_{i},4b_{0})}$
under the map $w\to\frac{w_{i}-w}{b_{0}}$.
Then for any $1\leq i\leq M$ and for $\sigma=\nu_{i},\tilde{\nu}_{i}$, we have
$\sigma(B(z,r))\leq e^{t/10}\delta^{-7\eta}r^{\alpha}\quad\text{for all $r\geq\delta_{0}/b_{0}$.}$
Thus, applying Lemma 2.2 with $\nu_{i},\tilde{\nu}_{i}$ (for any $1\leq i\leq M$), there exists a subset $E_{i}\subset B(w_{i},b_{0})$, with
$\nu_{i}(B(w_{i},b_{0})\setminus E_{i})\ll\delta^{\varepsilon_{0}}$
and for every $w\in E_{i}$, there is a subset $J_{w}$ with $|[0,1]\setminus J_{w}|\ll\delta^{\varepsilon_{0}}$ so that if $r\in J_{w}$, then for $\sigma=\frac{1}{\mu(B(w_{i},b_{0}))}\mu|_{B(w_{i},b_{0})}$ and $\sigma=\frac{1}{\mu(B(w_{i},4b_{0}))}\mu|_{B(w_{i},4b_{0})}$,
(2.5) $\sigma(\{w:|(1,r,\tfrac{r^{2}}{2})\cdot\tfrac{w-w^{\prime}}{b_{0}}|\leq e^{-t}\})$ $\ll e^{t/10}\delta^{-7\eta}e^{-49t/100}$
$\ll e^{-0.3t},$
where we used $e^{-t}\leq\delta^{100\eta}$ in the second inequality.
Let $F^{\prime\prime}_{\rm hm}=(F_{\rm hm}\cap\Xi_{\delta})\cap(\bigcup_{i}E_{i})$. Then
$\mu(F^{\prime\prime}_{\rm hm})\gg\delta^{\eta}$
For every $w\in F^{\prime\prime}_{\rm hm}$, let $\hat{I}(w)=I(w)\cap J_{w}$. Then $|\hat{I}(w)|\geq\frac{1}{2}\delta^{\eta}$. Moreover, for every $w\in F^{\prime\prime}_{\rm hm}$ and every $r\in\hat{I}(w)$,
$\{w^{\prime}\in B(w,b_{0}):\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq\delta\}\subset\\ \{w^{\prime}\in B(w_{i},4b_{0}):|(1,r,r^{2})\cdot\tfrac{w-w^{\prime}}{b_{0}}|\leq e^{-t}\}$
where $w\in B(w_{i},b_{0})$ (see (1.3) for the definition of $\pi_{r}=\pi_{t,r}$ and recall that $b_{0}=\delta^{1-3\eta}$). Applying (2.5) with $\sigma=\frac{1}{\mu(B(w_{i},4b_{0}))}\mu|_{B(w_{i},4b_{0})}$,
(2.6) $\mu(\{w^{\prime}\in B(w,b_{0}):\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq\delta\}\ll e^{-0.3t}\mu(B(w_{i},4b_{0}))$
$\ll C_{0}e^{-0.3t}b_{0}^{\alpha}<C_{0}e^{-0.2t}\delta^{\alpha-7\eta}$ ,
where in the last inequality, we used $e^{-t}\leq\delta^{100\eta}$.
In view of (2.6), for every $w\in F^{\prime\prime}_{\rm hm}$ and all $r\in\hat{I}(w)$, there exists some $b>b_{0}=\delta^{1-3\eta}$ so that (2.4) holds. That is:
(2.7) $m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\gg C_{0}e^{-t/10}\delta^{\alpha-17\eta}.$
Altogether, combining (2.3), (2.4), and (2.7), we have found a subset $F^{\prime\prime}_{\rm hm}\subset F_{\rm hm}$ with $\mu(F^{\prime\prime}_{\rm hm})\gg\delta^{\eta}$ and for every $w\in F^{\prime\prime}_{\rm hm}$ a subset $\hat{I}(w)\subset[\frac{1}{2},1]$ with $|\hat{I}(w)|\gg\delta^{\eta}$ so that the following holds. For every $w\in F^{\prime\prime}_{\rm hm}$ and all $w\in\hat{I}(w)$, there exists $b=b(w,r)\geq\delta^{1-3\eta}$ so that
$m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\gg C_{0}e^{-t/10}\delta^{\alpha-17\eta}$
Now applying pigeonhole principle (and Fubini’s theorem), there exists $b\geq\delta^{1-3\eta}$ and $F^{\prime}_{\rm hm}\subset F_{\rm hm}^{\prime\prime}$ with
$\mu(F^{\prime}_{\rm hm})\gg\delta^{3\eta/2}$
so that for all $w\in F^{\prime}_{\rm hm}$, we have
|$|\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-7\eta}\}|\geq\delta^{3\eta/2}.$
The proof is complete. ∎
Proof of Lemma 2.1.
Assuming $C$ is large enough, we may assume $\delta$ is small throughout the proof.
In view of Lemma 2.3, see in particular (2.1), we will replace $F_{\rm hm}$ by $F_{\rm hm}^{\prime}$ and assume that there is some $b\geq\delta^{1-3\eta}$ so that (2.2) holds for all $w\in F_{\rm hm}$.
For every $w\in F_{\rm hm}$, set
$I^{\prime}(w)=\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w)|\mathsf{D}_{b}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-7\eta}\}.$
Then $|I^{\prime}(w)|\gg\delta^{3/2}$, see (2.2). Choose three subsets $I^{\prime}_{j}(w)\subset\hat{I}^{\prime}(w)$ for $j=1,2,3$, which satisfy the following properties
(2.8) $|I^{\prime}_{j}(w)|\gg\delta^{2\eta}\qquad\text{and}\qquad{\rm dist}(I^{\prime}_{i}(w),I^{\prime}_{j}(w))\gg\delta^{3\eta/2}\quad\text{for $i\neq j$}$
For $j=1,2,3$, define
$E_{j}(w)=\{w^{\prime}\in F\cap\mathsf{D}_{b}(w):\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq\delta\text{ for some $r\in I^{\prime}_{j}(w)$}\}.$
We claim
(2.9) $\mu(E_{j}(w))\geq C_{0}be^{-t/10}\delta^{\alpha-1-6\eta}\qquad\text{for $j=1,2,3$.}$
Fix one $j$ and cover $I^{\prime}_{j}(w)$ with intervals $J_{1},\ldots,J_{N}$ of size $C^{\prime}\delta/b$ for some $C^{\prime}$ which will be chosen to be large. Thus
(2.10) $N\gg\delta^{2\eta}/(C^{\prime}\delta/b)=b\delta^{2\eta}/(C^{\prime}\delta)\gg b\delta^{2\eta-1}.$
Since $b\geq\delta^{1-3\eta}$, (2.10) in particular implies that $N\gg\delta^{-\eta}$.
For each $J_{i}$, let $r_{i}\in J_{i}\cap I^{\prime}_{j}(w)$. Discarding at most half of $J_{i}$’s, we will assume $|r_{i}-r_{i^{\prime}}|\geq C^{\prime}\delta/b$, and will continue to denote the collection by $J_{1},\ldots,J_{N}$. For every $1\leq i\leq N$, let
$E_{j,i}=\{w^{\prime}\in E_{j}(w):\|\pi_{r_{i}}(w)-\pi_{r_{i}}(w^{\prime})\|\leq\delta\}.$
Then $\mu(E_{j,i})\geq C_{0}e^{-\kappa t}\delta^{\alpha-7\eta}$. Moreover, $E_{j,i}\cap E_{j,i^{\prime}}=\emptyset$. Indeed if $w^{\prime}\in E_{j,i}\cap E_{j,i^{\prime}}$, then
$r_{i},r_{i^{\prime}}\in\{r\in[\tfrac{1}{2},1]:\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq\delta\}.$
Thus the set of $r\in[\tfrac{1}{2},1]$ so that
$|(1,r,\tfrac{r^{2}}{2})\cdot(w-w^{\prime})|\leq e^{-t}\delta\quad\text{and}$
$|(0,1,r)\cdot(w-w^{\prime})|\leq\delta$
has diameter $\geq C^{\prime}\delta/b$. Since $b\leq\|w-w^{\prime}\|\leq 2b$, we get a contradiction so long as $C^{\prime}$ is large enough.
Using (2.10) and $\mu(E_{j,i})\geq C_{0}e^{-\kappa t}\delta^{\alpha-7\eta}$, thus
$\mu(E_{j}(w))\gg N\cdot\mu(E_{j,i})\gg C_{0}be^{-t/10}\delta^{\alpha-1-5\eta},$
for $j=1,2,3$, as we claimed in (2.9).
Using $\mu(F_{\rm hm})\gg\delta^{3\eta/2}$ and (2.9), we conclude that
(2.11) $\mu(\{(w,w_{1},w_{2},w_{3})\in F_{\rm hm}\times F^{3}:w_{j}\in E_{j}(w)\})\gg\\ C_{0}^{3}b^{3}e^{-0.3t}\delta^{3\alpha-3-13\eta}$
We now find an upper bound for the measure of the set on the left side of (2.11). To that end, fix some $(w_{1},w_{2},w_{3})\in F^{3}$ so that there exists some $w\in F_{\rm hm}$ with $w_{j}\in E_{j}(w)$. In particular, we have $w_{j}\in\mathsf{D}_{b}(w)$, where $b\geq\delta^{1-3\eta}$, and
(2.12) $\|w_{1}-w_{j}\|\leq 4b,\qquad\text{for $j\in\{1,2,3\}$}.$
Let $r_{j}\in I^{\prime}_{j}(w)$ be so that $\|\pi_{r_{j}}(w)-\pi_{r_{j}}(w_{j})\|\leq\delta$. Then $|r_{i}-r_{j}|\geq\delta^{3\eta/2}$ for $i\neq j$ and
$(1,r_{j},r_{j}^{2})\cdot w=(1,r_{j},r_{j}^{2})\cdot w_{j}+O(e^{-t}\delta)\quad\text{for $j\in\{1,2,3\}$}.$
Thus $w$ belongs to a set with diameter $\ll e^{-t}\delta^{1-4.5\eta}$. This and (2.12) imply
$\mu(\{(w,w_{1},w_{2},w_{3})\in F_{\rm hm}\times F^{3}:w_{j}\in E_{j}(w)\})\ll C_{0}^{3}e^{-\alpha t}\delta^{\alpha-4.5\eta\alpha}b^{2\alpha}$
Comparing this upper bound and (2.11), we conclude
$b^{3}e^{-3t/10}\delta^{3\alpha-3-13\eta}\ll e^{-\alpha t}\delta^{\alpha-4.5\eta\alpha}b^{2\alpha},$
which implies
$\delta^{2\alpha-3-13\eta+4.5\eta\alpha}b^{3-2\alpha}\ll e^{(-\alpha+0.3)t}.$
Now using $b\geq\delta^{1-3\eta}$, $03-2\alpha>$ 0, and the above, we conclude that
(2.13) $\delta^{2\alpha-3-13\eta+4.5\eta\alpha}\delta^{(3-2\alpha)(1-3\eta)}=\delta^{-13\eta+4.5\eta\alpha-(9-6\alpha)\eta}\ll e^{(-\alpha+0.3)t}$
However, $-13\eta+4.5\eta\alpha-(9-6\alpha)\eta\leq-6\eta$, since $03-2\alpha>$ 0. Assuming $\delta$ is small enough and recalling that $3/10<\alpha$, (2.13) cannot hold. This contradiction completes the proof. ∎
Proof of Theorem 1.1
We now complete the proof of Theorem 1.1, which is based on Lemma 2.1 and Fubini’s theorem.
Recall that for every for all $r\in[\tfrac{1}{2},1]$ and all $w\in F$, we put
$m^{\delta}(\pi_{r}(w))=\mu\{w^{\prime}:\|\pi_{r}(w)-\pi_{r}(w^{\prime})\|\leq\delta\}$
Let $\varepsilon>0$ and let $\eta=\varepsilon/20$. For all $r\in[\frac{1}{2},1]$, let
$F_{\rm bad}(r)=\{w:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\}.$
The claim in Theorem 1.1 follows if we show that there is a subset $I_{\delta}\subset[\frac{1}{2},1]$ with $|[\frac{1}{2},1]\setminus I_{\delta}|\ll\delta^{\eta/2}$ so that for all $r\in I_{\delta}$, we have
$\mu(F_{\rm bad}(r))\ll\delta^{\eta/2}$
Let $C>1$, and assume that there exists a subset $I_{\rm bad}\subset[\frac{1}{2},1]$ with $|I_{\rm bad}|\geq C\delta^{\eta/2}$ so that for all $r\in I_{\rm bad}$, we have
$\mu(F_{\rm bad}(r))\geq C\delta^{\eta/2}.$
We will show this leads to a contradiction provided $C$ is large enough.
Equip $[\tfrac{1}{2},1]\times F$ with the product measure $\operatorname{Leb}\times\mu$. Let
$\mathsf{E}=\{(r,w)\in[\tfrac{1}{2},1]\times F:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\},$
and for every $w\in F$, let $\mathsf{E}_{w}=\{r:(r,w)\in\mathsf{E}\}$. The above then implies that
$\operatorname{Leb}\times\mu(\mathsf{E})\geq C^{2}\delta^{\eta}.$
Set $F^{\prime}=\{w\in F:|\mathsf{E}_{w}|\geq C\delta^{\eta}\}$. Then using Fubini’s theorem, we conclude
$\mu(F^{\prime})\geq\tfrac{1}{2}C^{2}\delta^{\eta}.$
Moreover, in in view of the definitions, for every $w\in F^{\prime}$
|$|\{r\in[\tfrac{1}{2},1]:m^{\delta}(\pi_{r}(w))\geq C_{0}e^{-t/10}\delta^{\alpha-18\eta}\}|\geq C\delta^{\eta}$
This contradicts Lemma 2.1 provided $C$ is large enough.
The proof of complete. ∎
3. Proof of Proposition 1.2
Let us now turn to the proof of Proposition 1.2. As it was already mentioned, we will prove this theorem using [GGW22, Thm. 2.1]. Let us begin by recalling the notation used in  [GGW22, Thm. 2.1] and Proposition 1.2.
Recall that for all $r\in[0,1]$, we put $\xi(r)=(\tfrac{r}{1!},\tfrac{r^{2}}{2!},\ldots,\tfrac{r^{n}}{n!})\subset\mathbb{R}^{n}$. For all $1\leq k\leq n$, let $\mathfrak{p}_{r}^{(k)}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ be the orthogonal projection onto the space spanned by $\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}$. That is:
$\mathfrak{p}_{r}^{(k)}(w)=\Bigl{(}w\cdot\xi^{\prime}(r),w\cdot\xi^{(2)}(r),\ldots,w\cdot\xi^{(k)}(r)\Bigr{)}$
where $w\cdot v$ is the usual inner product on $\mathbb{R}^{n}$.
3.1 Theorem ([GGW22]).
Let $1\leq k\leq n$ and let $0<\alpha\leq k$. Let $\mu$ be the uniform measure on a finite set $F\subset B_{\mathbb{R}^{n}}(0,1)$ satisfying
$\mu(B_{\mathfrak{r}}(w,\delta)\cap\Theta)\leq C_{0}\delta^{\alpha}\qquad\text{for all $w$ and all $\delta\geq\delta_{0}$}$
where $C_{0}>0$.
Let $0<\varepsilon<10^{-4}\alpha$. For every $\delta\geq\delta_{0}$, there exists a subset $J_{\delta}\subset[0,1]$ with $|[0,1]\setminus J_{\delta}|\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}$ so that the following holds. Let $r\in J_{\delta}$, then there exists a subset $F_{\delta,r}\subset F$ with
$\mu(F\setminus F_{\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}$
such that for all $w\in F_{\delta,r}$ we have
$\mu\Bigl{(}\{w^{\prime}\in F:\|\mathfrak{p}_{r}^{(k)}(w)-\mathfrak{p}_{r}^{(k)}(w^{\prime})\|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}\delta^{\alpha-\varepsilon}$
Proof.
We deduce this from [GGW22, Thm. 2.1]. The argument is more or less standard. Indeed it is similar to the deduction of Theorem 1.1 from Lemma 2.1 and to (a finitary version of) the argument in [GGW22, §2].
Since $1\leq k\leq n$ is fixed throughout the argument, we will denote $\mathfrak{p}_{r}^{(k)}$ by $\mathfrak{p}_{r}$. Adapting the notation $m^{\delta}$ from the previous section: for every $r\in[\frac{1}{2},1]$ and all $w\in F$, put
$m^{\delta}(\mathfrak{p}_{r}(w))=\mu\{w^{\prime}:\|\mathfrak{p}_{r}(w)-\mathfrak{p}_{r}(w^{\prime})\|\leq\delta\}.$
Let $D_{1},\ldots$ be large constants which will be explicated later. Let $\eta=\varepsilon/D_{1}$. For all $r\in[\frac{1}{2},1]$, let
$F_{\rm bad}(r)=\{w:m^{\delta}(\mathfrak{p}_{r}(w))\geq C_{0}\delta^{\alpha-D_{1}\eta}\}.$
Assume contrary to the claim in Theorem 3.1 that there exists a subset $I_{\rm bad}\subset[\frac{1}{2},1]$ with $|I_{\rm bad}|\geq D_{2}\delta^{\eta^{2}/2}$ so that for all $r\in I_{\rm bad}$, we have
$\mu(F_{\rm bad}(r))\geq D_{2}\delta^{\eta^{2}/2}.$
We will get a contradiction with [GGW22, Thm. 2.1], provided that $D_{i}$’s are large enough.
First note that, for every $r\in I_{\rm bad}$, the number of $\delta$-boxes $\{\mathsf{B}_{i,r}\}$ required to cover $\mathfrak{p}_{r}(F_{\rm bad}(r))$ is $\leq D_{3}C_{0}^{-1}\delta^{-\alpha+D_{1}\eta}$. Following [GGW22], let $\mathcal{T}_{r}=\{\mathbb{T}_{i,r}\}$ where $\mathbb{T}_{i,r}=\mathfrak{p}_{r}^{-1}(\mathsf{B}_{i,r})\cap B_{\mathbb{R}^{n}}(0,1)$; note that
(3.1) #$\#\mathcal{T}_{r}\leq D_{3}C_{0}^{-1}\delta^{-\alpha+D_{1}\eta}.$
Select a maximal $\delta$-separated subset $\Lambda_{\delta}\subset I_{\rm bad}$ and extend this to a maximal $\delta$-separated subset $\hat{\Lambda}_{\delta}$ of $[\frac{1}{2},1]$. The definition of $F_{\rm bad}$
Let $\rho$ denote the uniform measure on $\hat{\Lambda}_{\delta}$, and as it was done in the proof of Theorem 1.1, equip $\hat{\Lambda}_{\delta}\times F$ with the product measure $\rho\times\mu$, and let
$\mathsf{E}$ =$=\{(r,w)\in\Lambda_{\delta}\times F:m^{\delta}(\pi_{r}(w))\geq C_{0}\delta^{\alpha-D_{1}\eta}\}$
=$=\{(r,w)\in\Lambda_{\delta}\times F:w\in F_{\rm bad}(r)\}.$
Then the above implies that $\rho\times\mu(\mathsf{E})\geq D_{2}^{2}\delta^{\eta^{2}}$.
For every $w\in F$, let $\mathsf{E}_{w}=\{r\in\Lambda_{\delta}:(r,w)\in\mathsf{E}\}$, and set
$F^{\prime}=\{w\in F:\rho(\mathsf{E}_{w})\geq D_{2}\delta^{\eta^{2}}\}.$
Then using Fubini’s theorem, we conclude $\mu(F^{\prime})\geq\tfrac{1}{2}D_{2}^{2}\delta^{\eta^{2}}$.
Recall that $\rho$ is the normalized counting measure on $\Lambda_{\delta}$. The above definitions thus imply
$\sum_{r\in\Lambda_{\delta}}1_{\mathbb{T}_{r}}(w)\geq D_{3}\delta^{\eta^{2}-1}\qquad\text{for all $w\in F^{\prime}$}$
where $D_{3}=O(D_{2})$ and the implied constant is absolute. Let $\mu^{\prime}$ denote the restriction of $\mu$ to $F^{\prime}$. Now by [GGW22, Thm 2.1], applied with $\eta^{2}$, $\mu^{\prime}$ and $\{\mathcal{T}_{r}:r\in\Lambda_{\delta}\}$, we have
$\sum_{r\in\Lambda_{\delta}}\#\mathcal{T}_{r}$ $\geq D_{4}(n,\varepsilon,\alpha)C_{0}^{-1}\mu^{\prime}(\mathbb{R}^{n})\delta^{-1-\alpha+D\eta}$
$\geq\tfrac{1}{2}D_{2}^{2}D_{4}(n,\varepsilon,\alpha)C_{0}^{-1}\delta^{\eta^{2}}\delta^{-1-\alpha+D\eta}$
where $D=10^{10n}$ and in the second line we used $\mu^{\prime}(\mathbb{R}^{n})=\mu(F^{\prime})\geq\tfrac{1}{2}D_{2}^{2}\delta^{\eta^{2}}$. Thus there exists some $r\in\Lambda_{\delta}$ so that
(3.2) #$\#\mathcal{T}_{r}\geq\tfrac{1}{2}C_{0}^{-1}D_{4}(n,\varepsilon,\alpha)D_{2}^{2}\delta^{-\alpha+(D+1)\eta}.$
Now comparing (3.2) and (3.1) we get a contradiction so long as $D_{1}$ is large enough and $\delta$ is small enough. The proof is complete. ∎
We now turn to the proof of Proposition 1.2.
Proof of Proposition 1.2.
Let $\delta$ and $k$ be as in Proposition 1.2. As before, we let $\mu$ be the uniform measure on a finite set $F$.
Apply Theorem 3.1 with $\delta^{\prime}:=e^{-t}\delta\geq\delta_{0}$ and with $k$ and $k-1$. Then for every $r\in J_{\delta}$ and $w\in F_{\delta^{\prime},r}$, we have
(3.3) $\mu\Bigl{(}\{w^{\prime}\in F:\|\mathfrak{p}_{r}^{(\ell)}(w)-\mathfrak{p}_{r}^{(\ell)}(w^{\prime})\|\leq e^{-t}\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}(e^{-t}\delta)^{\min(\alpha,\ell)-\varepsilon}$
for $\ell=k,k-1$.
Also note that if $10<\alpha\leq k-$ 1, then
$\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\}\subset\\ \{w^{\prime}\in F:\|\mathfrak{p}_{r}^{(k-1)}(w)-\mathfrak{p}_{r}^{(k-1)}(w^{\prime})\|\leq e^{-t}\delta\}.$
Thus (3.3), with $\ell=k-1$ implies
(3.4) $\mu(\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\})\ll_{\varepsilon}C_{0}e^{-\alpha t}\delta^{\alpha-\varepsilon}$
as we claimed.
Assume now that $k-1<\alpha\leq k$. Let $Q\subset{\rm Span}\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}\simeq\mathbb{R}^{k}$ denote the box $\delta^{\prime}\times\cdots\times\delta^{\prime}\times\delta$ centered at the origin, where directions correspond directions of $\xi^{(i)}(r)$. That is: $Q$ is box of size $\delta^{\prime}$ in the direction of $\xi^{(i)}(r)$ for all $11\leq i\leq k-$ 1 and of size $\delta$ in the direction of $\xi^{(k)}(r)$. We cover
$Q^{\prime}:=(Q+\pi_{t,r}(w))\cap\pi_{t,r}(F_{\delta^{\prime},r})$
with $\ll e^{t}$ many boxes $B_{i}$ of size $\delta^{\prime}$. Then (3.3), applied with $2\delta^{\prime}$, implies
(3.5) $\mu(Q^{\prime})$ $\ll_{\varepsilon}(C_{0}(2\delta^{\prime})^{\alpha-\varepsilon})\cdot e^{t}$
$\ll_{\varepsilon}C_{0}e^{(1-\alpha)t}\delta^{\alpha-\varepsilon}.$
Finally, note that
$\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\}\subset Q+\pi_{t,r}(w).$
This and (3.5) complete the proof in this case as well. ∎

### v2

@@
 In this paper, we study a closely related family of projections. The consideration of these families are mainly motivated by the aforementioned applications to homogeneous dynamics.
-Let us fix some notation in order to state the main results of this paper. For every $t\geq 0$ and $r\in[0,1]$, let $\pi_{t,r}:\mathbb{R}^{3}\to\mathbb{R}^{2}$ be
+Indeed we first study the problem in dimension $3$, where an elementary argument inspired by [OV18] will be used. Then we use the main results in [GGW22] to study a similar problem in all dimensions.
+Let us fix some notation in order to state the results of this paper. For every $t\geq 0$ and $r\in[0,1]$, let $\pi_{t,r}:\mathbb{R}^{3}\to\mathbb{R}^{2}$ be
 (1.3) $\pi_{t,r}(x,y,z)=\Bigl{(}e^{t}(x+ry+\tfrac{r^{2}}{2}z),y+rz\Bigr{)}$
-The following is the main theorem in this paper.
+The following is one of the results of this paper.
 1.1 Theorem.
-Let $21\leq\alpha\leq 3/$ 2, and let $0<\delta_{0}\leq 1$. Let $F\subset B_{\mathbb{R}^{3}}(0,1)$ be a finite set satisfying the following:
-#$\#(B_{\mathbb{R}^{3}}(w,\delta)\cap F)\leq C_{0}\cdot\delta^{\alpha}\cdot(\#F)\quad\text{for all $w\in\mathbb{R}^{3}$ and all $\delta\geq\delta_{0}$}$
+Let $21\leq\alpha\leq 3/$ 2, and let $0<\delta_{0}\leq 1$. Let $F\subset B_{\mathbb{R}^{3}}(0,1)$ be a finite set and let $\mu$ denote the uniform measure on $F$. Assume that
+$\mu(B_{\mathbb{R}^{3}}(w,\delta))\leq C_{0}\cdot\delta^{\alpha}\quad\text{for all $w\in\mathbb{R}^{3}$ and all $\delta\geq\delta_{0}$}$
 where $C_{0}\geq 1$.
-Let $1000<\varepsilon<\alpha/$ 100. For every $\delta\geq e^{t}\delta_{0}$, there is a subset $I_{\delta}\subset[1/2,1]$ with
-$|[1/2,1]\setminus I_{\delta}|\ll_{\varepsilon}\delta^{\star\varepsilon}$
-so that the following holds. Let $r\in I_{\delta}$, then there exists $F_{\delta,r}\subset F$ with
-#$\#(F\setminus F_{\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon}\cdot(\#F)$
-such that for all $w\in F_{\delta,r}$, we have
-#($\#\Bigl{(}\{w^{\prime}\in F_{\delta,r}:\|\pi_{t,r}(w^{\prime})-\pi_{t,r}(w)|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}e^{-t/10}\delta^{\alpha-\varepsilon}\cdot(\#F)$
+Let $1000<\varepsilon<\alpha/$ 100.
+For every $\delta\geq e^{t}\delta_{0}$, there is a subset there exists $\mathcal{E}=\mathcal{E}_{{t},\delta}\subset[0,1]$ with $|\mathcal{E}|\ll_{\varepsilon}\delta^{\star\varepsilon}$ so that the following holds.
+Let $r\in[0,1]\setminus\mathcal{E}$, then there exists $F_{t,\delta,r}\subset F$ with
+$\mu(F\setminus F_{t,\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon}$
+such that for all $w\in F_{t,\delta,r}$, we have
+$\mu\Bigl{(}\{w^{\prime}\in F_{t,\delta,r}:\|\pi_{t,r}(w^{\prime})-\pi_{t,r}(w)|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}e^{-t/10}\delta^{\alpha-\varepsilon}$
 Remark.
@@
 Also, for a Borel subset $\mathsf{B}\subset\mathbb{R}^{d}$, we denote the Lebesgue measure of $\mathsf{B}$ by $|\mathsf{B}|$.
-We also state a similar result in all dimensions $n\geq 3$, which relies on the aforementioned work of [GGW22].
-Let $12\leq k\leq n-$ 1. For all $t\geq 0$ and all $r\in[0,1]$, let $\pi_{t,r}^{(k)}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ be
-$\pi_{t,r}^{(k)}(w)=\Bigl{(}e^{t}w\cdot\xi^{\prime}(r),e^{t}w\cdot\xi^{\prime\prime}(r),\ldots,e^{t}w\cdot\xi^{(k-1)}(r),w\cdot\xi^{(k)}(r)\Bigr{)}$
-where $w\cdot v$ is the usual inner product on $\mathbb{R}^{n}$ and
-$\xi(r)=(\tfrac{r}{1!},\tfrac{r^{2}}{2!},\ldots,\tfrac{r^{n}}{n!})\subset\mathbb{R}^{n}.$
-Using [GGW22, Thm. 2.1] and some elementary arguments, one can establish the following
-1.2 Proposition.
-Let $0<\alpha\leq k$, and let $0<\delta_{0}\leq 1$. Let $F\subset B_{\mathbb{R}^{n}}(0,1)$ be a finite set satisfying the following:
-#$\#(B_{\mathbb{R}^{n}}(w,\delta)\cap F)\leq C_{0}\cdot\delta^{\alpha}\cdot(\#F)\quad\text{for all $w\in F$ and all $\delta\geq\delta_{0}$}$
-where $C_{0}\geq 1$.
-Let $1000<\varepsilon<\alpha/$ 100. For every $\delta\geq e^{t}\delta_{0}$, there is a subset $I_{\delta}\subset[1/2,1]$ with
-$|[1/2,1]\setminus I_{\delta}|\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}$
-so that the following holds. Let $r\in I_{\delta}$, then there exists $F_{\delta,r}\subset F$ with
-#$\#(F\setminus F_{\delta,r})\ll_{\varepsilon}\delta^{\star\varepsilon^{2}}\cdot(\#F)$
-such that for all $w\in F_{\delta,r}$, we have
-#($\#\Bigl{(}\{w^{\prime}\in F_{\delta,r}:\|\pi_{t,r}(w^{\prime})-\pi_{t,r}(w)|\leq\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}e^{-\alpha^{\prime}t}\delta^{\alpha-\varepsilon}\cdot(\#F)$
-where $\alpha^{\prime}=\alpha$ if $10<\alpha\leq k-$ 1 and $\alpha^{\prime}=\alpha-1$ otherwise.
-It is worth noting that Proposition 1.2 does not imply Theorem 1.1; indeed $\alpha^{\prime}=\alpha-1<1/10$ for $101\leq\alpha<11/$ 10.
-That said, it seems plausible that the approach developed in [GGW22] can be modified to prove a stronger version of Proposition 1.2 which then would imply (and perhaps improve) Theorem 1.1.
-However, our proof of Theorem 1.1 relies on much softer arguments and is an adaptation of [OV18].
-Also as previously mentioned, our precise formulations here (which at least at first glance may appear somewhat ad hoc) are motivated by recent applications in homogeneous dynamics [LM23, LMW22, LMWY23].
+As mentioned earlier, the proof of Theorem 1.1 is based on elementary arguments inspired by [OV18].
+However, this approach does not readily extend to higher dimensions, nor does it provide a clear path to overcoming the barrier $\alpha\leq 3/2$ in the context of Theorem 1.1.
+We now present a similar general result on dimension improvement in irreducible ${\rm{SL}}_{2}(\mathbb{R})$-representations.
+In dimensions $3,5,7$, and $11$ results of this nature drive the Margulis function estimate in [LMWY25], see also [LMW22].
+The proof relies on deep ingredients.
+In particular, our argument makes crucial use of [GGW22, Thm.
+2.1].
+Let $V$ be an irreducible representation of ${\rm{SL}}_{2}(\mathbb{R})$ with $\dim V=n+1$.
+We fix a norm on $V$.
+Note that since all norm on a finite dimensional vector space are equivalent, different choice of norms will only create a constant factor in the dimension estimate.
+Let
+$u_{r}=\begin{pmatrix}1&r\\ &1\end{pmatrix}$
+and
+$a_{t}=\begin{pmatrix}e^{{t}/2}&\\ &e^{-{t}/2}\end{pmatrix}.$
+We have the following result on dimension improvement under the push by $a_{t}u_{r}$ for a discretized measure in an irreducible representation of ${\rm{SL}}_{2}(\mathbb{R})$.
+1.2 Theorem.
+Let $\alpha\in(0,n+1)$, $C\geq 1$ and $0<\delta_{0}\leq 1$. Let $F$ be a finite set in $B_{V}(0,1)$ and let $\mu$ denote the uniform measure on $F$. Assume that
+$\mu(B_{V}(x,\delta))\leq C\delta^{\alpha}$
+for all $\delta\in[\delta_{0},1]$. Then the following holds for all $\epsilon\in(0,\frac{1}{10^{4}n}\min\{\alpha,n+1-\alpha\})$.
+For all $\delta\in[\delta_{0},\frac{1}{100}]$ and ${t}\gg_{\epsilon}1$ with $e^{\frac{n}{2}{t}}\delta_{0}\leq\delta\leq e^{-\frac{n}{2}{t}}$, there exists $\mathcal{E}=\mathcal{E}_{{t},\delta}\subset[0,1]$ with $|\mathcal{E}|\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ satisfying the following.
+For all $r\in[0,1]\setminus\mathcal{E}$, there exists $F_{t,\delta,r}\subset F$ with $\mu(F\setminus F_{t,\delta,r})\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ so that for all $w\in F_{t,\delta,r}$, we have
+$\mu(\{w^{\prime}\in F_{t,\delta,r}:\|a_{t}u_{r}(w^{\prime})-a_{t}u_{r}(w)\|\leq\delta\})\ll_{\epsilon}Ce^{-\frac{\varpi(\alpha){t}}{2}}\delta^{\alpha-\epsilon}$
+where
+$\varpi(\alpha)=\max\{n\alpha-\lfloor\alpha\rfloor(\lfloor\alpha\rfloor+1),\lfloor\alpha\rfloor(2n-\lfloor\alpha\rfloor+1)-n\alpha\}.$
+All the implicit constant here depends only on $n$.
+Remark.
+Note that if $\alpha$ varies in compact subset of $(0,n+1)$, $\varpi(\alpha)$ is uniformly bounded away from 00.
+Remark.
+When $n=2$, the representation is the adjoint representation of ${\rm{SL}}_{2}(\mathbb{R})$. We have
+$\frac{1}{2}\varpi(\alpha)=\begin{cases}\alpha&0<\alpha\leq 1,\\ \max\{2-\alpha,\alpha-1\}&1<\alpha\leq 2,\\ 3-\alpha&2<\alpha\leq 3.\end{cases}$
+See Figure 1.
+33331111
+Figure 1. Graph of the piecewise function $\varpi(\cdot)$
+Remark.
+We remark that Theorem 1.2 is a generalization of Theorem 1.1 when ${t}=|\log\delta|$.
+In this case, the pre-image of a $\delta$-disk under the expanded projections in Theorem 1.1 is of size $\delta^{2}\times\delta\times 1$, which coincide with the pre-image of a $\delta$-ball under $a_{t}u_{r}$.
 2. Proof of Theorem 1.1
 In this section we will prove Theorem 1.1. Let us begin by fixing some notation which will be used throughout this section.
-Let $\mu$ denote the normalized counting measure on $F$:
+Recall that $\mu$ denotes the uniform measure on $F$. That is:
 $\mu(\mathsf{B})=\frac{\#(\mathsf{B}\cap F)}{\#F}\qquad\text{for any Borel set $\mathsf{B}\subset\mathbb{R}^{3}$}.$
@@
 The proof of complete. ∎
-3. Proof of Proposition 1.2
-Let us now turn to the proof of Proposition 1.2.
-As it was already mentioned, we will prove this theorem using [GGW22, Thm.
-2.1].
-Let us begin by recalling the notation used in  [GGW22, Thm.
-2.1] and Proposition 1.2.
-Recall that for all $r\in[0,1]$, we put $\xi(r)=(\tfrac{r}{1!},\tfrac{r^{2}}{2!},\ldots,\tfrac{r^{n}}{n!})\subset\mathbb{R}^{n}$.
-For all $1\leq k\leq n$, let $\mathfrak{p}_{r}^{(k)}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ be the orthogonal projection onto the space spanned by $\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}$.
-That is:
+3. Proof of Theorem 1.2
+We will prove Theorem 1.2 in this section. Before starting the proof let us fix some notation.
+For all $r\in[0,1]$, we put $\xi(r)=(\tfrac{r}{1!},\tfrac{r^{2}}{2!},\ldots,\tfrac{r^{n}}{n!},\tfrac{r^{n+1}}{(n+1)!})\in\mathbb{R}^{n+1}$. For all $11\leq k\leq n+$ 1, let
+$\mathfrak{p}_{r}^{(k)}:\mathbb{R}^{n+1}\to\mathbb{R}^{k}$
+be the projection onto the space spanned by $\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}$ defined by
 $\mathfrak{p}_{r}^{(k)}(w)=\Bigl{(}w\cdot\xi^{\prime}(r),w\cdot\xi^{(2)}(r),\ldots,w\cdot\xi^{(k)}(r)\Bigr{)}$
-where $w\cdot v$ is the usual inner product on $\mathbb{R}^{n}$.
+where $w\cdot v$ is the usual inner product on $\mathbb{R}^{n+1}$. We define $\mathfrak{p}^{(0)}_{r}$ to be the zero map for all $r\in[0,1]$.
+This family of projections has the following two properties.
+  1. (1)
+    The projection $\mathfrak{p}^{(n+1)}_{r}$ is exactly the action of $u_{r}$ in this representation.
+  2. (2)
+    The difference between $\mathfrak{p}_{r}^{(k)}$ and the orthogonal projection to the space spanned by $\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}$ is a bi-Lipschitz map where the Lipschitz constant depends only on $n$.
+Recall that we let
+$u_{r}=\begin{pmatrix}1&r\\ &1\end{pmatrix}$
+and
+$a_{t}=\begin{pmatrix}e^{{t}/2}&\\ &e^{-{t}/2}\end{pmatrix}.$
+With this normalization, there exists a basis for the irreducible representation $V$ so that we can identify $V\cong\mathbb{R}^{n+1}$ and for all $w\in\mathbb{R}^{n+1}$,
+$u_{r}.w={}$ $\biggl{(}w\cdot\xi^{(1)}(r),\cdots,w\cdot\xi^{(n)}(r)\biggr{)},$
+$a_{{t}}.w={}$ $(e^{\frac{n}{2}{t}}w_{1},e^{\frac{n-2}{2}{t}}w_{2}\cdots,e^{-\frac{n}{2}{t}}w_{n+1}).$
+For a probability measure $\mu$ supported on $B_{V}(0,1)$, we define the following notions. If $\mu(A)>0$, we let
+$\mu_{A}=\tfrac{1}{\mu(A)}\mu|_{A}$
+be the normalized restriction of $\mu$ on $A$.
+In this section, we will call cubes of the form $\prod_{i=1}^{n+1}[\frac{n_{i}}{2^{k}},\frac{n_{i}+1}{2^{k}})$ dyadic cubes.
+For a dyadic cube $Q\subset\mathbb{R}^{n+1}$, we set ${\rm{Hom}}_{Q}$ to be the homothety that maps $Q$ to $[0,1)^{n+1}$.
+We set
+$\mu^{Q}=({\rm{Hom}}_{Q})_{\ast}\mu_{Q}$
+to be the rescaled normalized restriction of $\mu$ on $Q$.
+We also refer to $\mu^{Q}$ as the conditional measure of $\mu$ on $Q$.
+We use the notion $\mathcal{D}_{\rho}$ for the collection of dyadic $\rho$-cubes and $A_{Q}=A\cap Q$.
+As it was already mentioned, we will prove Theorem 1.2 using [GGW22, Thm. 2.1]. We record the following consequence of  [GGW22, Thm. 2.1].
 3.1 Theorem ([GGW22]).
-Let $1\leq k\leq n$ and let $0<\alpha\leq k$. Let $\mu$ be the uniform measure on a finite set $F\subset B_{\mathbb{R}^{n}}(0,1)$ satisfying
-$\mu(B_{\mathfrak{r}}(w,\delta)\cap\Theta)\leq C_{0}\delta^{\alpha}\qquad\text{for all $w$ and all $\delta\geq\delta_{0}$}$
+Let $11\leq k\leq n+$ 1 and let $0<\alpha\leq k$. Let $\mu$ be the uniform measure on a finite set $F\subset B_{\mathbb{R}^{n+1}}(0,1)$ satisfying
+$\mu(B_{\mathbb{R}^{n+1}}(w,\delta))\leq C_{0}\delta^{\alpha}\qquad\text{for all $w$ and all $\delta\geq\delta_{0}$}$
 where $C_{0}>0$.
@@
 Indeed it is similar to the deduction of Theorem 1.1 from Lemma 2.1 and to (a finitary version of) the argument in [GGW22, §2].
-Since $1\leq k\leq n$ is fixed throughout the argument, we will denote $\mathfrak{p}_{r}^{(k)}$ by $\mathfrak{p}_{r}$.
+See also [JL24, Section 2].
+Since $11\leq k\leq n+$ 1 is fixed throughout the argument, we will denote $\mathfrak{p}_{r}^{(k)}$ by $\mathfrak{p}_{r}$.
 Adapting the notation $m^{\delta}$ from the previous section: for every $r\in[\frac{1}{2},1]$ and all $w\in F$, put
@@
 First note that, for every $r\in I_{\rm bad}$, the number of $\delta$-boxes $\{\mathsf{B}_{i,r}\}$ required to cover $\mathfrak{p}_{r}(F_{\rm bad}(r))$ is $\leq D_{3}C_{0}^{-1}\delta^{-\alpha+D_{1}\eta}$.
-Following [GGW22], let $\mathcal{T}_{r}=\{\mathbb{T}_{i,r}\}$ where $\mathbb{T}_{i,r}=\mathfrak{p}_{r}^{-1}(\mathsf{B}_{i,r})\cap B_{\mathbb{R}^{n}}(0,1)$; note that
+Following [GGW22], let $\mathcal{T}_{r}=\{\mathbb{T}_{i,r}\}$ where $\mathbb{T}_{i,r}=\mathfrak{p}_{r}^{-1}(\mathsf{B}_{i,r})\cap B_{\mathbb{R}^{n+1}}(0,1)$; note that
 (3.1) #$\#\mathcal{T}_{r}\leq D_{3}C_{0}^{-1}\delta^{-\alpha+D_{1}\eta}.$
 Select a maximal $\delta$-separated subset $\Lambda_{\delta}\subset I_{\rm bad}$ and extend this to a maximal $\delta$-separated subset $\hat{\Lambda}_{\delta}$ of $[\frac{1}{2},1]$.
-The definition of $F_{\rm bad}$
 Let $\rho$ denote the uniform measure on $\hat{\Lambda}_{\delta}$, and as it was done in the proof of Theorem 1.1, equip $\hat{\Lambda}_{\delta}\times F$ with the product measure $\rho\times\mu$, and let
@@
 Now by [GGW22, Thm 2.1], applied with $\eta^{2}$, $\mu^{\prime}$ and $\{\mathcal{T}_{r}:r\in\Lambda_{\delta}\}$, we have
-$\sum_{r\in\Lambda_{\delta}}\#\mathcal{T}_{r}$ $\geq D_{4}(n,\varepsilon,\alpha)C_{0}^{-1}\mu^{\prime}(\mathbb{R}^{n})\delta^{-1-\alpha+D\eta}$
+$\sum_{r\in\Lambda_{\delta}}\#\mathcal{T}_{r}$ $\geq D_{4}(n,\varepsilon,\alpha)C_{0}^{-1}\mu^{\prime}(\mathbb{R}^{n+1})\delta^{-1-\alpha+D\eta}$
 $\geq\tfrac{1}{2}D_{2}^{2}D_{4}(n,\varepsilon,\alpha)C_{0}^{-1}\delta^{\eta^{2}}\delta^{-1-\alpha+D\eta}$
-where $D=10^{10n}$ and in the second line we used $\mu^{\prime}(\mathbb{R}^{n})=\mu(F^{\prime})\geq\tfrac{1}{2}D_{2}^{2}\delta^{\eta^{2}}$. Thus there exists some $r\in\Lambda_{\delta}$ so that
+where $D=10^{10(n+1)}$ and in the second line we used $\mu^{\prime}(\mathbb{R}^{n+1})=\mu(F^{\prime})\geq\tfrac{1}{2}D_{2}^{2}\delta^{\eta^{2}}$. Thus there exists some $r\in\Lambda_{\delta}$ so that
 (3.2) #$\#\mathcal{T}_{r}\geq\tfrac{1}{2}C_{0}^{-1}D_{4}(n,\varepsilon,\alpha)D_{2}^{2}\delta^{-\alpha+(D+1)\eta}.$
 Now comparing (3.2) and (3.1) we get a contradiction so long as $D_{1}$ is large enough and $\delta$ is small enough. The proof is complete. ∎
-We now turn to the proof of Proposition 1.2.
-Proof of Proposition 1.2.
-Let $\delta$ and $k$ be as in Proposition 1.2. As before, we let $\mu$ be the uniform measure on a finite set $F$.
-Apply Theorem 3.1 with $\delta^{\prime}:=e^{-t}\delta\geq\delta_{0}$ and with $k$ and $k-1$. Then for every $r\in J_{\delta}$ and $w\in F_{\delta^{\prime},r}$, we have
-(3.3) $\mu\Bigl{(}\{w^{\prime}\in F:\|\mathfrak{p}_{r}^{(\ell)}(w)-\mathfrak{p}_{r}^{(\ell)}(w^{\prime})\|\leq e^{-t}\delta\}\Bigr{)}\ll_{\varepsilon}C_{0}(e^{-t}\delta)^{\min(\alpha,\ell)-\varepsilon}$
-for $\ell=k,k-1$.
-Also note that if $10<\alpha\leq k-$ 1, then
-$\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\}\subset\\ \{w^{\prime}\in F:\|\mathfrak{p}_{r}^{(k-1)}(w)-\mathfrak{p}_{r}^{(k-1)}(w^{\prime})\|\leq e^{-t}\delta\}.$
-Thus (3.3), with $\ell=k-1$ implies
-(3.4) $\mu(\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\})\ll_{\varepsilon}C_{0}e^{-\alpha t}\delta^{\alpha-\varepsilon}$
-as we claimed.
-Assume now that $k-1<\alpha\leq k$.
-Let $Q\subset{\rm Span}\{\xi^{\prime}(r),\ldots,\xi^{(k)}(r)\}\simeq\mathbb{R}^{k}$ denote the box $\delta^{\prime}\times\cdots\times\delta^{\prime}\times\delta$ centered at the origin, where directions correspond directions of $\xi^{(i)}(r)$.
-That is: $Q$ is box of size $\delta^{\prime}$ in the direction of $\xi^{(i)}(r)$ for all $11\leq i\leq k-$ 1 and of size $\delta$ in the direction of $\xi^{(k)}(r)$.
-We cover
-$Q^{\prime}:=(Q+\pi_{t,r}(w))\cap\pi_{t,r}(F_{\delta^{\prime},r})$
-with $\ll e^{t}$ many boxes $B_{i}$ of size $\delta^{\prime}$. Then (3.3), applied with $2\delta^{\prime}$, implies
-(3.5) $\mu(Q^{\prime})$ $\ll_{\varepsilon}(C_{0}(2\delta^{\prime})^{\alpha-\varepsilon})\cdot e^{t}$
-$\ll_{\varepsilon}C_{0}e^{(1-\alpha)t}\delta^{\alpha-\varepsilon}.$
-Finally, note that
-$\{w^{\prime}:\|\pi_{t,r}(w)-\pi_{t,r}(w^{\prime})\|\leq\delta\}\subset Q+\pi_{t,r}(w).$
-This and (3.5) complete the proof in this case as well. ∎
+Now we prove Theorem 1.2.
+Proof.
+Without loss of generality, we assume both $e^{\frac{n{t}}{2}}\delta$ and $\delta$ are dyadic numbers.
+This will only result a constant factor depending only on the ambient space in the final estimate.
+For simplicity, let $\rho=e^{\frac{n{t}}{2}}\delta$ and $k=\lfloor\alpha\rfloor$. Fix $\delta$ satisfying $e^{\frac{n}{2}{t}}\delta_{0}\leq\delta\leq e^{-\frac{n}{2}{t}}$.
+Let $P$ be the tube of size
+$e^{-n{t}}\times\cdots\times e^{-{t}}\times 1$
+Centered at the origin, with its sides aligned along the coordinate axes, via the identification $V\cong\mathbb{R}^{n+1}$.
+We note that $u_{r}^{-1}a_{{t}}^{-1}B(x,\delta)$ is a translate of $u_{r}^{-1}.P$.
+We denote it by $w+u_{r}^{-1}.P$.
+We will prove the theorem by estimating $\mu(w+u_{r}^{-1}.P)$.
+First note that since tubes $w+u_{r}^{-1}.P$ lies in $O(1)$ many dyadic $\rho$-cubes, it suffices to estimate $\mu_{Q}(w+u_{r}^{-1}.P)$ for those dyadic $\rho$-cubes $Q$ intersecting $w+u_{r}^{-1}.P$ non-trivially.
+To this end, we first study the dimension condition for those conditional measure $\mu^{Q}$.
+For all $s\geq\rho^{-1}\delta_{0}$, we have
+$\mu^{Q}(B(x,s))={}$ $\mu_{Q}({\rm{Hom}}_{Q}^{-1}B(x,s))$
+[…]
+[…]
+Now we study the relation between $\mu_{Q}(w+u_{r}^{-1}.P)$ with the projections $\Bigl{\{}\mathfrak{p}_{r}^{(l)}\Bigr{\}}_{r,l}$.
+This follows from the standard ‘projection and slicing’ picture, as we now explicate .
+Namely, in our case, for all $l=0,\cdots,n+1$, there exists some rectangle $R_{l}(w)\subset\mathbb{R}^{l}$ with sides parallel to the axis and side-length
+$e^{-n{t}}\times\cdots\times e^{-(n-l+1){t}}$
+so that
+$w+u_{r}^{-1}.P\subseteq\Bigl{(}\mathfrak{p}_{r}^{(k)}\Bigr{)}^{-1}(R_{l}(w)).$
+It suffices to study $\Bigl{(}\mathfrak{p}_{r}^{(k)}\Bigr{)}_{\ast}\mu^{Q}(R_{l}(w))$. We will focus on the cases where $l=k$ and $l=k+1$.
+Apply Theorem 3.1 with the measures $\{\mu^{Q}\}_{Q}$, $\mathfrak{p}^{(k)}_{r}$, scale $e^{-n{t}}$, and $\epsilon/n$.
+Using Fubini’s theorem, there exists $\mathcal{E}_{{t},\delta}\subset[0,1]$ with $|[0,1]\setminus\mathcal{E}_{{t},\delta}|\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ so that for all $r\in[0,1]\setminus\mathcal{E}_{{t},\delta}$, there exists $\mathcal{D}_{\rho}(r)$ with
+$\sum_{Q\notin\mathcal{D}_{\rho}(r)}\mu(Q)\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}.$
+Moreover, for all $Q\in\mathcal{D}_{\rho}(r)$, there exists $F_{Q}(r)\subset F_{Q}$ with $\mu_{Q}(F_{Q}\setminus F_{Q}(r))\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ so that for all $w\in V$, we have
+$\mu^{Q}\Bigl{(}F_{Q}(r)\cap(w+u_{r}^{-1}.P)\Bigr{)}\leq{}$ $\Bigl{(}\mathfrak{p}_{r}^{(k)}\Bigr{)}_{\ast}\mu^{Q}(R_{k}(w))$
+$\ll_{\epsilon}{}$ $\frac{C\rho^{\alpha}}{\mu(Q)}\prod_{j=0}^{k-1}e^{-(n-j){t}}e^{\epsilon{t}}.$
+The last inequality follows from the fact that we can cover $R_{k}(w)$ with $O(\prod_{j=0}^{k-1}e^{j{t}})$ many $e^{-n{t}}$ balls.
+Let $F_{r}^{(k)}=\sqcup_{Q\in\mathcal{D}_{\rho}(r)}F_{Q}(r)$, we have
+[…]
+Now we apply Theorem 3.1 with $\{\mu^{Q}\}_{Q}$, $\mathfrak{p}^{(k+1)}_{r}$, scale $e^{-n{t}}$, and $\epsilon/n$.
+Using Fubini’s theorem, again, there exists $\mathcal{E}_{{t},\delta}^{\prime}\subset[0,1]$ with $|[0,1]\setminus\mathcal{E}_{{t},\delta}|\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ so that for all $r\in[0,1]\setminus\mathcal{E}_{{t},\delta}^{\prime}$, there exists $\mathcal{D}_{\rho}(r)^{\prime}$ with
+$\sum_{Q\notin\mathcal{D}_{\rho}(r)^{\prime}}\mu(Q)\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}.$
+Moreover, for all $Q\in\mathcal{D}_{\rho}(r)^{\prime}$, there exists $F_{Q}(r)^{\prime}\subset F_{Q}$ with $\mu_{Q}(F_{Q}\setminus F_{Q}(r)^{\prime})\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$ so that for all $w\in V$, we have
+$\mu^{Q}(F_{Q}(r)^{\prime}\cap(w+u_{r}^{-1}.P))\leq{}$ $(\mathfrak{p}_{r}^{(k+1)})_{\ast}\mu^{Q}(R_{k+1}(w))$
+$\ll_{\epsilon}{}$ $\frac{C\rho^{\alpha}}{\mu(Q)}\Biggl{(}\prod_{j=0}^{k}e^{j{t}}\Biggr{)}e^{-n\alpha{t}}e^{\epsilon{t}}.$
+As before, the last inequality follows from the fact that we can cover $R_{k+1}(w)$ via $O(\prod_{j=0}^{k}e^{j{t}})$ many $e^{-n{t}}$ balls.
+Let $F_{r}^{(k+1)}=\sqcup_{Q\in\mathcal{D}_{\rho}(r)}F_{Q}(r)$, we have
+[…]
+Let $\mathcal{E}=\mathcal{E}_{{t},\delta}\cup\mathcal{E}_{{t},\delta}^{\prime}$ and for all $r\notin\mathcal{E}$, let $F_{t,\delta,r}=F_{r}^{(k)}\cap F_{r}^{(k+1)}$. We have
+$|[0,1]\setminus\mathcal{E}|\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}$
+and
+$\mu(F\setminus F_{r})\ll_{\epsilon}e^{-\star\epsilon^{2}{t}}.$
+Combining (3.3) and (3.4), the theorem follows. ∎